OpenStudy (anonymous):

Solve and check. 7 3/4 = x/8 A. 1/4 B. 62 C. 1/62 D. 4

OpenStudy (petiteme):

B. :)

OpenStudy (charlotte123):

@studyislandhelp it is B :D 7 3/4 = 31/4. 31/4 = x/8 *You can see that the denominator is multiplied by 2, so u multiply the numerator by 2 :D 31*2 = 62 :) So, x = 62 :D Meaning 31/4 = 62/8 :)